A firm produces output Q by using capital K and labor L in fixed proportions, i.e. Q = F (K ,L ) = min {K, L/3}. The price of a unit of labor is w = 6, the price of a unit of capital is r = 2 and the price of output is p = 20.

a) Draw the isoquant for Q = 8.

b) Find the marginal product of labor.

Suppose that (in part c and d) the firm's capital is fixed at K = 8.

c) Write down the firm's short-run profit as a function of output. Does the firm make a profit or loss and how much when it produces 8 units of output?

d) What is the optimal quantity of labor the firm would use?
Answer #1

8) a) a= min (K, 43), W=6, 8=2, P=20, Q=8, 80 8= Min (Kg 4/3) at Eqm 80 K-8, & L=24 . K=43 b) Now. If 43 <K, then Q = 43 Then